# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ B.4 Gradient Flow of Invalid Loss
In this section, we show that the penalty for invalid token samples through the arg max function on predicted logits can be effectively back-propagated through the model parameters via our softmax scaling strategy that bypasses the undifferentiable arg max function. Here, we will denote the predicted probability for the token $k=\arg\max_{j}\big{(}\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\big{)}$ with the highest probability as $x_{\theta,k}^{(\ell)}$ and all remaining token probabilities as $x_{\theta,j}^{(\ell)}$ for $j=[1\dots|\mathcal{V}|]$.
@@
Conversely, the parameters of the remaining tokens $x^{(\ell)}_{\theta,j}$ are updated such that the predicted probability of sampling the other tokens increases in proportion to their original softmax probabilities. This prevents extreme changes in the predicted probabilities of the remaining tokens and ensures that the token distribution remains relatively consistent with the previous iteration.
@@
Here, we show the theoretical basis of the gradient update that effectively updates parameters to reduce the position-specific token probabilities that result in invalid sequence samplings and push the model predictions toward other high-likelihood tokens.
@@ C.1 RoFormer Architecture Details
C.2 Target-Binding Prediction Model
@@
C.3 Boosted Trees for Peptide SMILES Property Prediction
@@
C.4 PDB Docking Structures
Table 11: PDB structures used for docking.
Protein PDB
GFAP 6A9P
TfR 3KAS
GLP-1R 3C5T
AMHR2 7L0J
GLAST 5LM4
NCAM1 2HAZ
RBX1 1LDJ
@@ C.5 Evaluation Metrics
Validity is defined as the fraction of peptide SMILES that pass our SMILES2PEPTIDE filter, indicating that it translates to a synthesizable peptide.
@@ D.1 Case Study for Time-Dependent Multi-Objective Guidance
Some properties of peptides require more intense guidance towards specific structural or sub-structural features, while others may only require small changes in the sidechain composition or non-natural modifications. To enable the prioritization of properties during guidance, we introduce a time-dependent multi-objective guidance strategy that guides the generation based on only a subset of properties depending on the current iteration number of the MCTS search (Algorithm 2). To achieve this, we define a $K$-dimensional vector $\mathbf{i}=[i_{1},i_{2},\dots,i_{K}]$ where each $i_{k}$ is the iteration number to begin guidance for the $k$th objective. Properties where $i_{k}=0$ are used to guide all iterations, whereas properties where $i_{k}>1$ are used to guide only the iterations from $i_{k}\to N_{\text{iter}}$.
@@
Our time-dependent guidance operates as follows. During the expansion and rollout steps on iteration $i$, the rolled-out child sequences $\mathbf{x}_{s,i}$ that are non-dominated across the sub-vector of property scores $\mathbf{s}_{i}=[s_{k}\;|\;i_{k}\leq i\leq N_{\text{iter}}]$ dependent on the iteration $i$ are added to the Pareto-optimal set $\mathcal{P}^{*}$. Therefore, $\mathbf{x}_{s}$ does not need to be non-dominant in the properties $k$ where $i_{k}>i$. Similarly, only the sequences $\tilde{\mathbf{x}}\in\mathcal{P}^{*}$ that are dominated in the subset of properties represented in $\mathbf{s}_{i}(\tilde{\mathbf{x}})$ are removed from $\mathcal{P}^{*}$.
$\mathcal{P}^{\prime*}$ $=\mathcal{P}^{*}\cup\big{\{}(\mathbf{z}_{s},\mathbf{s}(\mathbf{x}_{s}))\;|\;\forall\tilde{\mathbf{x}}\in\mathcal{P}^{*}\;\;\mathbf{s}_{i}(\mathbf{x}_{s})\succeq\mathbf{s}_{i}(\tilde{\mathbf{x}})\big{\}}$ (92)
@@
$\mathcal{P}^{\prime*}$ $=\mathcal{P}^{*}\setminus\big{\{}\tilde{\mathbf{x}}\;|\;\exists\mathbf{x}_{s}\in\text{children}(\mathbf{z}_{t})\;\text{s.t.}\;\mathbf{s}_{i}(\mathbf{x}_{s})\succ\mathbf{s}_{i}(\tilde{\mathbf{x}})\big{\}}$ (93)
@@
Then, during selection, we only consider the cumulative rewards $\mathbf{U}_{i}=[u_{k}\;|\;i_{k}\leq i\leq N_{\text{iter}}]$ for the properties where $i_{k}$ s.t. $i_{k}\leq i\leq N_{\text{iter}}$ to form the set of Pareto-optimal children nodes that will be chosen from uniformly at random.
$\mathcal{P}^{\prime*}_{\text{select}}$ $=\mathcal{P}^{*}_{\text{select}}\cup\big{\{}\mathbf{z}_{s}\;|\;\forall\tilde{\mathbf{x}}\in\mathcal{P}^{*}\;\;\mathbf{U}_{i}(\mathbf{x}_{s})\succeq\mathbf{U}_{i}(\tilde{\mathbf{x}})\big{\}}$ (94)
@@
$\mathcal{P}^{\prime*}$ $=\mathcal{P}^{*}\setminus\big{\{}\tilde{\mathbf{x}}\;|\;\exists\mathbf{x}_{s}\in\text{children}(\mathbf{z}_{t})\;\text{s.t.}\;\mathbf{U}_{i}(\mathbf{x}_{s})\succ\mathbf{U}_{i}(\tilde{\mathbf{x}})\big{\}}$ (95)
Then, we select the next node $\mathbf{z}_{s}\sim\mathcal{P}^{\prime*}_{\text{select}}$ similarly to the time-independent case.
@@
To test this strategy, we generated 100 peptides conditioned only on membrane permeability for the first 50 iterations since we found it as the most challenging property to optimize. Then we conditioned all properties including membrane permeability, binding affinity to GFAP, solubility, hemolysis, and non-fouling. We show that during the first 50 iterations, all properties except membrane permeability show relatively constant average scores, whereas the permeability score increased (Fig LABEL:). Then, after the 50 iteration mark, GFAP binding affinity and solubility curves increased significantly while the hemolysis and non-fouling curves increased slightly for the remainder of the iterations (Fig LABEL:). Although all the results in this paper leverage peptides without time-dependent guidance, this serves as a proof of concept for future experiments varying the start times across properties and testing the generated peptides to refine additional properties at later time steps where the generated sequences are already constrained to certain predefined substructures.
@@ D.2 Ablation Studies
In this section, we report the performance of our multi-objective MCTS-guided discrete diffusion model over different hyperparameter settings. Specifically, we discuss the effects of varying the number of iterations, the number of expanded children nodes, the weight of invalid penalty, the weight of binding affinity reward, and the batched sampling method.
@@
Number of Iterations. The number of iterations $N_{\text{iter}}$ determines the number of selection, expansion, rollout, backpropagation loops are executed in a single MCTS run. In addition, $N_{\text{iter}}$ is the maximum value of $t$ or the number of unmasking steps that can be executed before the rollout begins, which corresponds to the maximum tree depth. We found that equating the number of diffusion steps $T$ to the number of MCTS iterations $N_{\text{iter}}$ results in convergence on the property prediction scores as the selection process becomes biased towards a single unmasking scheme. As shown in Figure 6, all property scores converge for $N_{\text{iter}}=T=128$, which we use for the remainder of the study.
Tokenization Scheme. To evaluate the effect of different tokenization methods on the generation quality, we experimented with three different tokenization schemes: pre-trained PeptideCLM SMILES Pair Encoding (SPE) tokenization, and Atom Pair Encoding (APE) tokenization trained on the 30K fine-tuning dataset of valid peptide SMILES.
@@ Appendix E Algorithms
Algorithm 1 outlines the forward pass diffusion process (training) for PepMDLM, our unconditional peptide SMILES generator. Algorithms 2, 3, 4, and 5 describe PepTune, our MCTS-guided peptide SMILES generator. Algorithms 6 and 7 describe the bond mask function and peptide sequence decoder which can also act as a validity filter.
@@
Output: Pre-trained unconditional MDLM for peptide SMILES generation $\mathbf{x}_{\theta}(\mathbf{z}_{t})$
@@
     Sample $\mathbf{z}_{t}\sim\text{Cat}(\mathbf{z}_{t};(\alpha_{t}(\mathbf{x}_{0}))\mathbf{x}_{0}+(1-\alpha_{t}(\mathbf{x}_{0})\mathbf{m})$
@@
     $\mathcal{L}^{\infty}_{\text{NELBO}}\leftarrow\frac{1}{|B|}\sum_{\mathbf{z}_{t}\in B}\bigg{(}\frac{1}{t}\sum_{\ell:\mathbf{x}^{\ell}\neq\mathbf{b}}-\log\langle\mathbf{x}_{\theta}(\mathbf{z}^{\ell}_{t}),\mathbf{x}^{\ell}\rangle+\frac{w}{t}\sum_{\ell:\mathbf{x}^{\ell}=\mathbf{b}}-\log\langle\mathbf{x}_{\theta}(\mathbf{z}^{\ell}_{t}),\mathbf{x}^{\ell}\rangle\bigg{)}$
@@
Inputs: Pre-trained MDLM denoiser $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$, score function $\mathbf{s}(\mathbf{x}):\mathcal{V}^{L}\to\mathbf{R}^{K}$ containing classifiers for $K$ objectives $s_{1},s_{2},\dots,s_{K}$, number of time steps $T$, number of iterations $N_{\text{iter}}$
@@
         $\text{children}(\mathbf{z}_{t})\leftarrow$ BatchedReverseStep$(\mathbf{z}_{t})$
@@
              $\mathbf{z}_{t}\leftarrow\mathbf{z}_{s,i}$
@@
                  $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})\leftarrow$ReverseDiffusionStep$(\mathbf{z}_{t})$
                  $\mathbf{z}_{s}\sim p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$
@@
              $\mathbf{x}_{s,i}\leftarrow\arg\max\big{(}\mathbf{x}_{\theta}(\mathbf{z}_{t},1)\big{)}$$\triangleright$ get clean sequence
              $\mathbf{s}(\mathbf{x}_{s,i})\leftarrow\mathbf{s}(\mathbf{x}_{s,i})$$\triangleright$ compute score vector
              $\mathbf{r}(\mathbf{z}_{s,i}),\mathcal{P}^{*}\leftarrow$UpdateParetoFront$\big{(}\mathcal{P}^{*},(\mathbf{z}_{s,i},\mathbf{s}(\mathbf{z}_{s,i}))\big{)}$$\triangleright$ add sequence if non-dominant
@@
              $\mathbf{U}(\mathbf{z})\leftarrow\mathbf{U}(\mathbf{z})+\mathbf{r}$
@@
              $\mathbf{z}\leftarrow\text{parent}(\mathbf{z})$
@@
     $\mathbf{u}_{s}\sim\text{Uniform}(0,1)^{L\times|\mathcal{V}|}$
     $\mathbf{G}_{s}\leftarrow-\log\left(-\log(\mathbf{u}_{s}+\epsilon)+\epsilon\right)$
@@
              $\tilde{p}_{\theta,i}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t})\leftarrow\log p_{\theta}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t}\big{)}+\mathbf{G}_{s,i}$$\triangleright$ batched Gumbel-max distributions
@@
              $\tilde{p}_{\theta,i}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t})\leftarrow\text{SM}\bigg{(}\text{topk}\big{\{}\log(p_{\theta}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t})\big{)}+\mathbf{G}_{s,i}\big{\}}\bigg{)}$$\triangleright$ batched top $k$ sampling
@@
     end for
     for $i=1\dots M$ do
         $\mathbf{z}_{s,i}\sim\tilde{p}_{\theta,i}$
@@
                       $\mathbf{U}(\mathbf{z}_{s,i})\leftarrow\frac{\mathbf{W}(\mathbf{z}_{s,i})}{N_{\text{visit}}(\mathbf{z}_{s,i})}$$\triangleright$ compute normalized cumulative reward vector
                       $\mathcal{P}^{*}_{\text{select}}\leftarrow$UpdateParetoFront$(\mathcal{P}^{*}_{\text{select}},(\mathbf{z}_{s,i},\mathbf{U}(\mathbf{z}_{s,i})))$$\triangleright$ update Pareto frontier

### v2

@@ B.4 Gradient Flow of Invalid Loss
In this section, we show that the penalty for invalid token samples through the argmax function on predicted logits can be effectively backpropagated through the model parameters via our softmax scaling strategy. Here, we will denote the predicted probability for the token $k=\arg\max_{j}\big{(}\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\big{)}$ with the highest probability as $x_{\theta,k}^{(\ell)}$ and all remaining token probabilities as $x_{\theta,j}^{(\ell)}$ for $j=[1\dots|\mathcal{V}|]$.
@@
In contrast, the parameters of the remaining tokens $x^{(\ell)}_{\theta,j}$ are updated so that the predicted probability of sampling the other tokens increases proportionally to their original softmax probabilities. This prevents extreme changes in the predicted probabilities of the remaining tokens and ensures that the token distribution remains relatively consistent with the previous iteration.
@@
Here, we show our invalidity loss effectively updates parameters to reduce the position-specific token probabilities that result in invalid sequence samplings and push the model predictions toward other high-likelihood tokens.
@@ C.1 RoFormer Architecture Details
C.2 PDB Docking Structures
Table 9: PDB structures used for docking.
Protein PDB
GFAP 6A9P
TfR 3KAS
GLP-1R 3C5T
AMHR2 7L0J
GLAST 5LM4
NCAM1 2HAZ
RBX1 1LDJ
C.3 Target-Binding Prediction Model
@@
C.4 Boosted Trees for Peptide SMILES Property Prediction
@@ C.5 Evaluation Metrics
Validity is defined as the fraction of peptide SMILES that pass our SMILES2PEPTIDE filter (Algorithm 8), indicating that it translates to a synthesizable peptide.
@@ D.1 Case Study for Time-Dependent Multi-Objective Guidance
Some properties of peptides require more intense guidance towards specific structural or sub-structural features, while others may only require small changes in the side chain composition or non-natural modifications. To enable the prioritization of properties during guidance, we introduce a time-dependent multi-objective guidance strategy that guides the generation based on only a subset of properties depending on the current iteration number of the MCTS search. To achieve this, we define a $K$-dimensional vector $\mathbf{i}=[i_{1},i_{2},\dots,i_{K}]$ where each $i_{k}$ is the iteration number to begin guidance for the $k$th objective. Properties where $i_{k}=0$ are used to guide all iterations, whereas properties where $i_{k}>1$ are used to guide only the iterations from $i_{k}\to N_{\text{iter}}$.
@@
Our time-dependent guidance operates as follows. During the expansion and rollout steps on iteration $i$, the rolled-out child sequences $\mathbf{x}_{s,i}$ that are non-dominated across the sub-vector of property scores $\mathbf{s}_{i}=[s_{k}\;|\;i_{k}\leq i\leq N_{\text{iter}}]$ dependent on the iteration $i$ is added to the Pareto-optimal set $\mathcal{P}^{*}$. Therefore, $\mathbf{x}_{s}$ does not need to be non-dominated in the properties $k$ where $i_{k}>i$. Similarly, only the sequences $\mathbf{x}^{*}\in\mathcal{P}^{*}$ that become dominated when adding $\mathbf{x}_{s}$ in the subset of properties represented in $\mathbf{s}_{i}(\mathbf{x}^{*})$ are removed from $\mathcal{P}^{*}$.
$\mathcal{P}^{\prime*}$ $=\mathcal{P}^{*}\cup\big{\{}(\mathbf{z}_{s},\mathbf{s}(\mathbf{x}_{s}))\;|\;\forall\mathbf{x}^{*}\in\mathcal{P}^{*}\;\;\mathbf{s}_{i}(\mathbf{x}_{s})\succeq\mathbf{s}_{i}(\mathbf{x}^{*})\big{\}}$ (89)
@@
$\mathcal{P}^{\prime*}$ $=\mathcal{P}^{*}\setminus\big{\{}\mathbf{x}^{*}\;|\;\exists\mathbf{x}_{s}\in\text{children}(\mathbf{z}_{t})\;\text{s.t.}\;\mathbf{s}_{i}(\mathbf{x}_{s})\succ\mathbf{s}_{i}(\mathbf{x}^{*})\big{\}}$ (90)
@@
Then, during selection, we only consider the cumulative rewards $\mathbf{W}_{i}=[W_{k}\;|\;i_{k}\leq i\leq N_{\text{iter}}]$ for the properties where $i_{k}$ s.t. $i_{k}\leq i\leq N_{\text{iter}}$ when computing the selection score vector $\mathbf{U}_{i}$ to form the Pareto-optimal selection set.
$\mathcal{P}^{\prime*}_{\text{select}}$ $=\mathcal{P}^{*}_{\text{select}}\cup\big{\{}\mathbf{z}_{s}\;|\;\forall\mathbf{x}^{*}\in\mathcal{P}^{*}_{\text{select}}\;\;\mathbf{U}_{i}(\mathbf{z}_{t},\mathbf{z}_{s})\succeq\mathbf{U}_{i}(\mathbf{z}_{t},\mathbf{z}^{*})\big{\}}$ (91)
@@
$\mathcal{P}^{\prime*}_{\text{select}}$ $=\mathcal{P}^{*}_{\text{select}}\setminus\big{\{}\mathbf{z}^{*}\;|\;\exists\mathbf{z}_{s}\in\text{children}(\mathbf{z}_{t})\;\text{s.t.}\;\mathbf{U}_{i}(\mathbf{z}_{t},\mathbf{z}_{s})\succ\mathbf{U}_{i}(\mathbf{z}_{t},\mathbf{z}^{*})\big{\}}$ (92)
Finally, we select the next node $\mathbf{z}_{s}\sim\mathcal{P}^{\prime*}_{\text{select}}$ uniformly at random from the Pareto-optimal selection set.
@@
To test this strategy, we generated 100 peptides conditioned only on membrane permeability for the first 50 iterations since we found it as the most challenging property to optimize. Then we conditioned all properties including membrane permeability, binding affinity to GFAP, solubility, hemolysis, and non-fouling. We show that during the first 50 iterations, all properties except membrane permeability show relatively constant average scores, whereas the permeability score increased (Figure 16). Then, after the 50 iteration mark, GFAP binding affinity and solubility curves increased significantly while the hemolysis and non-fouling curves increased slightly for the remainder of the iterations (Figure 16). Although all the results in this paper leverage peptides without time-dependent guidance, this serves as a proof of concept for future experiments varying the start times across properties to refine certain properties at later time steps where the generated sequences are already constrained to specific predefined substructures.
@@ D.2 Ablation Studies
In this section, we report the performance of our multi-objective MCTS-guided discrete diffusion model over different hyperparameter settings. Specifically, we discuss the effects of changing the number of expanded children nodes, the number of iterations, and the tokenization scheme.
@@
Number of Iterations. The number of iterations $N_{\text{iter}}$ determines the number of selection, expansion, rollout, and backpropagation loops executed in a single MCTS run. In addition, $N_{\text{iter}}$ is the maximum value of $t$ or the number of unmasking steps that can be executed before the rollout begins, which corresponds to the maximum tree depth. We found that equating the number of diffusion steps $T$ to the number of MCTS iterations $N_{\text{iter}}$ results in convergence on the property prediction scores as the selection process becomes biased towards a single unmasking scheme. As shown in Figure 6, all property scores converge for $N_{\text{iter}}=T=128$, which we use for the remainder of the study.
Tokenization Scheme. To evaluate the effect of different tokenization methods on the generation quality, we experimented with three different tokenization schemes: SMILES Pair Encoding (SPE) tokenization with the trained vocabulary used by PeptideCLM and Atom Pair Encoding (APE) tokenization for SMILES and SELFIES [97] representations. Overall, we found that the SPE tokenization scheme decreased perplexity and maintained precision while capturing common peptide motifs like bonds and recurring side chains.
@@ Appendix E Algorithms
Algorithm 1 outlines the forward pass diffusion process (training) for PepMDLM, our unconditional peptide SMILES generator. Algorithms 2, 3, 4, 5 and 6 describe PepTune, our MCTS-guided peptide SMILES generator. Algorithms 7 and 8 describe the bond mask function and peptide sequence decoder which can also act as a validity filter.
@@
Output: Trained unconditional MDLM for peptide SMILES generation $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$
@@
     Sample $\mathbf{z}_{t}\sim\text{Cat}(\mathbf{z}_{t};(\alpha_{t}(\mathbf{x}_{0}))\mathbf{x}_{0}+(1-\alpha_{t}(\mathbf{x}_{0}))\mathbf{m})$
@@
     $\mathcal{L}^{\infty}_{\text{NELBO}}\leftarrow\frac{1}{|B|}\sum_{\mathbf{x}_{0}\in B}\bigg{(}-\sum_{\ell:\mathbf{x}^{(\ell)}_{0}=\mathbf{b}}\frac{w}{t}\log\langle\mathbf{x}^{(\ell)}_{0},\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\rangle-\sum_{\ell:\mathbf{x}^{(\ell)}_{0}\neq\mathbf{b}}\frac{1}{t}\log\langle\mathbf{x}^{(\ell)}_{0},\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\rangle\bigg{)}$
@@
Algorithm 2 Reverse Diffusion Step
Inputs: Partially unmasked sequence at time $t$ $\mathbf{z}_{t}$
Output: Token probability distribution $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$ for all positions in the sequence with the state-dependent reverse posterior and SUBS parametrization
procedure ReverseDiffusionStep
     $s\leftarrow t-\frac{1}{T}$
     if $\mathbf{z}_{t}=\mathbf{m}$ then
     else if $\mathbf{z}_{t}\neq\mathbf{m}$ then
     end if
     return $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$
end procedure
@@
Inputs: Trained MDLM denoiser $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$, score function $\mathbf{s}(\mathbf{x}):\mathcal{V}^{L}\to\mathbb{R}^{K}$ containing classifiers for $K$ objectives $s_{1},s_{2},\dots,s_{K}$, number of time steps $T$, number of iterations $N_{\text{iter}}$
@@
         $\text{children}(\mathbf{z}_{t(n)})\leftarrow$ BatchedReverseStep$(\mathbf{z}_{t})$
@@
              $\mathbf{z}_{t(n)}\leftarrow\mathbf{z}_{s,i}$
@@
                  ${p}_{\theta,i}(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})\leftarrow$ReverseDiffusionStep$(\mathbf{z}_{t(n)})$
                  $\mathbf{z}_{t(n)}\leftarrow\arg\max{p}_{\theta,i}(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})$
@@
              $\mathbf{x}_{s,i}\leftarrow\arg\max{p}_{\theta,i}(\mathbf{x}_{s,i}|\mathbf{z}_{t(1)})$$\triangleright$ get clean sequence
[…]
              $\mathbf{r}(\mathbf{z}_{s,i}),\mathcal{P}^{*}\leftarrow$UpdateParetoFront$\big{(}\mathcal{P}^{*},(\mathbf{z}_{s,i},\mathbf{s}(\mathbf{z}_{s,i}))\big{)}$
@@
              $\mathbf{W}(\mathbf{z})\leftarrow\mathbf{W}(\mathbf{z})+\mathbf{r}$
@@
              $\mathbf{z}\leftarrow\text{parent}(\mathbf{z})$$\triangleright$ repeat for parent node until root node
@@
     for $\ell=1\dots L;i=1\dots M;j=1\dots|\mathcal{V}|$ do
         $u^{(\ell)}_{i,j}\sim\text{Uniform}(0,1)$
         $G^{(\ell)}_{i,j}\leftarrow-\log\left(-\log(u^{(\ell)}_{i,j}+\epsilon)+\epsilon\right)$
     end for
@@
              $\tilde{p}_{\theta,i}\big{(}\mathbf{z}_{s,i}|\mathbf{z}_{t})\leftarrow\log p_{\theta}\big{(}\mathbf{z}_{s,i}|\mathbf{z}_{t}\big{)}+\mathbf{G}_{i}$$\triangleright$ batched Gumbel-max distributions
              $\mathbf{z}_{s,i}\sim\tilde{p}_{\theta,i}\big{(}\mathbf{z}_{s,i}|\mathbf{z}_{t})$
@@
              $\tilde{p}_{\theta,i}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t})\leftarrow\text{SM}\bigg{(}\text{top}k\big{\{}\log p_{\theta}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t}\big{)}+\mathbf{G}_{i}^{(\ell)}\big{\}}\bigg{)}$$\triangleright$ batched top $k$ sampling
@@
         $\mathbf{z}_{s,i}\sim\tilde{p}_{\theta,i}\big{(}\mathbf{z}_{s,i}|\mathbf{z}_{t})$
@@
                       $\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i})\leftarrow\frac{\mathbf{W}(\mathbf{z}_{s,i})}{N_{\text{visit}}(\mathbf{z}_{s,i})}+c\cdot p_{\theta}(\mathbf{z}_{s,i}|\mathbf{z}_{t})\frac{\sqrt{N_{\text{visit}}(\mathbf{z}_{t})}}{1+N_{\text{visit}}(\mathbf{z}_{s,i})}$
                       $\mathcal{P}^{*}_{\text{select}}\leftarrow$UpdateParetoFront$\bigg{(}\mathcal{P}^{*}_{\text{select}},(\mathbf{z}_{s,i},\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i}))\bigg{)}$